Finding operating and free cash flows Consider the balance sheets and selected data from the income statement of Keith Corpor$1,610 2,740 December 31 2015 2014 $1,490 $1,040 1,790 1.250 2,010 1.850 2.920 2.820 $8,210 $6.960 $29,450 $28,060 14,750 13.
Answer #1

a) NOPAT = EBIT * (1 - taxes)

= 2740 * (1 - 0.4)

= $1644

b)

operating cash flow = NOPAT + Depreciation

= 1644 + 1610

= $3254

  c)

free cash flows = operating cash flow - change in fixed assets(investment in fixed assets) - change in net working capital

net working capital= current assets -   current liabilities

change in fixed assets = 29450 - 28060

= 1390

change in current assets(except inventory) = 8210 - 6960

= 1250

change in current liabilities = (1570+190 -1530 - 260)

  = -30

change in net working capital 1250 - (-30) = 1280

so free cash flow = 3254 - 1390 - 1280

= $584

d)

A - NOPAT = 1644

depreciation = 1610

both are approximately same size

since OCF = NOPAT + Depreciation

OCF is twice the size of NOPAT So this statement is true

C - this is also true because FCF arrives after deducting investment in fixed assets and changes in NWC

D - it can be clearly seen that kieth has postive cash flows from operating activities.so this statement is true